11 years ago, Hye was four times as old as her brother. In 1 year she will be twice as old as he is now. What are their ages now?
Answer by checkley71(8403)

11 YEARS AGE HYE=4X & HER BROTHER=X.
12 YEARS LATER HYE=(4X+12) & HER BROTHER=(X+12).
SO (4X+12)=2(X+12)
4X+12=2X+24
4X-2X=24-12
2X=12
X=6 WHICH IS HER BROTHERS AGE 11 YEARS AGO.
HYE'S AGE 11 YEARS AGO=4*6=24.
PROOF:
24+12=2(6+12)
36=2*18
36=36